OpenStudy (anonymous):

Quadrilateral ABCD has coordinates (3, 5), (5, 2), (11, 6), (9, 9). Quadrilateral ABCD is a: A: square because its length and width are both Square Root 10 units and adjacent sides are perpendicular. B: rhombus because its length and width are both Square Root 10 units and adjacent sides are not perpendicular. C: rectangle because its length is 2 Square Root 10 units, its width is Square root 10 units, and adjacent sides are perpendicular. D: trapezoid because it has only one pair of parallel sides.

OpenStudy (anonymous):

Distances: AB = SQRT(9+4) = SQRT(13) BC = SQRT(16+36) = SQRT(52) CD = SQRT(4+9) = SQRT(13) AD = SQRT(36+16) = SQRT(52) So, it must be a rectangle.

jhonyy9 (jhonyy9):

- so i think this can be rhombus

OpenStudy (anonymous):

Rhombus = equal sides.